Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company uses a standard cost system for all of its products. According to the standards that have been set for the seat covers, the factory should work 2,850 hours each month to produce 1,900 sets of covers. The standard costs associated with this level of production are:

Total

Per Set
of Covers

Direct materials

$

42,560

$

22.40

Direct labor

$

51,300

27.00

Variable manufacturing overhead (based on direct labor-hours)

$

6,840

3.60

$

53.00

During August, the factory worked only 2,800 direct labor-hours and produced 2,000 sets of covers. The following actual costs were recorded during the month:

Total

Per Set
of Covers

Direct materials (12,000 yards)

$

45,600

$

22.80

Direct labor

$

49,000

24.50

Variable manufacturing overhead

$

7,000

3.50

$

50.80

At standard, each set of covers should require 5.6 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

1

Material price variance

2400

F

Material price quantity

3200

U

2

Labor rate variance

1400

F

Labor efficiency variance

???

F

3

Variable overhead rate variance

280

U

Variable overhead efficiency variancy

480

F
Answer #1
1. Standard price of materials per yard = Standard direct materials per set of covers / Standard quantity of materials per set of covers = 22.40 / 5.6 4
Actual materials rate per yard = Actual total materials cost / Actual materials quantity = 45600 / 12000 3.8
Materials price variance = ( Actual price - Standard price ) * Actual quantity = ( 3.8 - 4 ) * 12000 2400 F
Standard quantity of materials for 2000 sets of covers = 2000 * Standard quantity of materials per set of covers = 2000 * 5.6 11200
Materials quantity variance = ( Actual quantity - Standard quantity ) * Standard price = ( 12000 - 11200 ) * 4 3200 U
2.
Standard labor hours per sets of cover = Standard hours for 1900 sets of covers / 1900 = 2850 / 1900 1.5
Standard rate of labor = Standard direct labor per set of cover / Standard labor hours per sets of covers = 27 / 1.5 18
Actual rate of labor = Actual total direct labor cost / Actual direct labor hours worked = 49000 / 2800 17.50
Labor rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 17.50 - 18 ) * 2800 1400 F
Standard direct labor hours for 2000 sets of covers = 2000 * Standard labor hours per sets of covers = 2000 * 1.5 3000
labor efficiency varaince = ( Actual hours - Standard hours ) * Standard rate = ( 2800 - 3000 ) * 18 3600 F
3.
Standard rate of variable overhead = Standard cost of variable overhead per sets of cover / Standard labor hours per sets of cover = 3.60 / 1.5 2.4
Actual rate of variable overhead = Actual total variable overhead cost / Actual direct labor hours worked = 7000 / 2800 2.5
Variable overhead rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 2.5 - 2.4 ) * 2800 280 U
Variable overhead efficiency variance = ( Actual hours - Standard hours ) * Standard rate = ( 2800 - 3000 ) * 2.4 480 F